You discover an annual invasive plant (which has discrete reproduction and dies in one year) in a previously un-infested ditch near your house. Upon further investigation, you determine that there are exactly 500 individuals in the population. If this population has a per capita birth rate of 20 plants per year, what is λ for this population (include 3 decimal places)?
Answer #1

= 500 (initial population)

=

where, B is number of births and D is number of deaths.

  is populaytion after time t ( here 1 year)

= 500 + (500 x 20) - (500) [ As per capita birth rate given we mulltiply it with initial population]

= 500 + 10000 - 500

= 10000 (population after 1 year)

Therefore,

Lambda =

= 10000/500

= 20.